OpenStudy (anonymous):

Find the perimeter of the shape:

OpenStudy (anonymous):

I have no idea how to do this, any help is greatly appreciated.

OpenStudy (rational):

perimeter = length around the shape

OpenStudy (rational):

how many sides do u see in the given shape ?

OpenStudy (anonymous):

4.

OpenStudy (rational):

yes ! find the length of each side and add them

OpenStudy (rational):

maybe start by finding the length of side \(AB\)

OpenStudy (anonymous):

I have no idea haha is it 3.5?

OpenStudy (rational):

Notice that AB is goind 4 grids in x, and 1 grid in y : so length of AB = \(\sqrt{4^2 + 1^2} = \sqrt{17}\)

OpenStudy (anonymous):

So 4.1? :O

OpenStudy (rational):

if u find all the lengths and add them you should get perimeter = 21.44 http://www.wolframalpha.com/input/?i=sqrt%284%5E2%2B1%5E2%29+%2B+sqrt%282%5E2%2B2%5E2%29+%2B+sqrt%282%5E2%2B7%5E2%29+%2B+sqrt%284%5E2%2B6%5E2%29

OpenStudy (anonymous):

Thank you, I'm not fully understanding perimeter I will have to study more.

OpenStudy (rational):

just practice few problems on pythagorean theorem / distance formula in `xy plane`